OpenStudy (khally92):

Set up only an exact iterated integral expression representing the area of the plane region bounded by

OpenStudy (khally92):

\[y^2=x+4~~~~and ~~~~~~x-y=2\]

OpenStudy (khally92):

using the order dydx ONLY

zepdrix (zepdrix):

Let's find the intersection of these functions :)

zepdrix (zepdrix):

\[\large\rm y^2-4=x\qquad\qquad\qquad x=y+2\]\[\large\rm y^2-4=y+2\]Solve for y to find the intersections. I let you do that easy step :) Let's call those values y1 and y2.

zepdrix (zepdrix):